If the following jobs are sequenced according to the SLACK rule then the maximum job tardiness (in days) is (assume zero for today's date)

Job

Processing Time (days)

Due Date

A

8

12

B

6

15

C

11

17

D

7

10

E

3

8
Answer #1

A F 1 Job B с D Processing Time (days) Due Date Slack 7 10 E Flow time Tradiness 7 0 15 3 2 D 3 3 A 3 12 4 4 E 3 8 5 18 10 5

Maximum job tardiness is 20 days

FOrmula

A B C D F 1 Job Processing Time (days) Due Date 2 D 7 10 3 A 8 12 Slack =C2-B2 =C3-B3 =C4-B4 =C5-B5 =C6-B6 E Flow time -SUM(B